# Evaluating Self-Supervised Learning in Medical Imaging: A Benchmark for Robustness, Generalizability, and Multi-Domain Impact

### v1

@@ Abstract
Self-supervised learning (SSL) has emerged as a promising paradigm in medical imaging, addressing the chronic challenge of limited labeled data in healthcare settings. While SSL has shown impressive results, existing studies in the medical domain are often limited in scope, focusing on specific datasets or modalities, or evaluating only isolated aspects of model performance. This fragmented evaluation approach poses a significant challenge, as models deployed in critical medical settings must not only achieve high accuracy but also demonstrate robust performance and generalizability across diverse datasets and varying conditions. To address this gap, we present a comprehensive evaluation of SSL methods within the medical domain, with a particular focus on robustness and generalizability. Using the MedMNIST dataset collection as a standardized benchmark, we evaluate 8 major SSL methods across 11 different medical datasets. Our study provides an in-depth analysis of model performance in both in-domain scenarios and the detection of out-of-distribution (OOD) samples, while exploring the effect of various initialization strategies, model architectures, and multi-domain pre-training. We further assess the generalizability of SSL methods through cross-dataset evaluations and the in-domain performance with varying label proportions (1%, 10%, and 100%) to simulate real-world scenarios with limited supervision. We hope this comprehensive benchmark helps practitioners and researchers make more informed decisions when applying SSL methods to medical applications.
@@ 1 Introduction
Medical image annotation is a resource-intensive task that requires specialized domain knowledge, making it significantly more costly and laborious than annotating natural images [79, 45, 63, 39]. The scarcity of labeled medical data, combined with the complexity of the annotation process, presents a significant challenge for building effective machine learning models in healthcare. Self-supervised learning has emerged as a powerful solution to these limitations, enabling models to learn rich representations from unlabeled data prior to task-specific fine-tuning with minimal labeled samples [24, 21, 40, 6, 2, 71, 38].
@@
While self-supervised learning (SSL) has proven effective in improving classification performance with limited labels [14, 29, 80, 13, 15], there has been limited work in evaluating the robustness and generalizability of learned models. For safe deployment in healthcare, models must perform reliably across diverse settings and recognize when to withhold predictions on out-of-distribution (OOD) samples. Although SSL methods have been evaluated for OOD detection in natural image contexts [48, 43, 36, 57, 31], their effectiveness in medical imaging remains largely under-researched. Berger et al. [9] explore OOD detection using supervised approaches in a clinical context, but focus solely on chest X-rays, limiting insights across other domains. Bozorgtabar et al. [11] evaluate SSL methods for OOD detection under domain shift but they only investigate two SSL approaches in a single medical domain. Although Cai et al. [12] provide a comprehensive study of SSL for anomaly detection, they focus specifically on anomaly detection SSL methods rather than general-purpose SSL approaches.
@@
Beyond OOD detection, an effective model should generalize across different tasks and modalities—a key requirement in medical imaging where datasets and conditions vary widely. A model that performs well across different datasets or modalities ensures continued diagnostic support even when some imaging modalities are inaccessible, reduces the need for extensive retraining, and ensures robust performance in diverse clinical settings. Azizi et al. [7] propose an SSL-based representation learning method and assess its robustness across various medical imaging tasks and domains. However, they do not compare different SSL methods, leaving a gap in research on the cross-domain generalizability of common SSL approaches.
@@
To address these gaps, our study introduces a comprehensive benchmark for evaluating the robustness and generalizability of SSL methods in medical imaging. We assess various SSL methods for OOD detection, comparing the effectiveness of convolutional networks (ResNet-50) with transformer-based architectures (ViT-Small). We further assess the generalizability of learned representations by training linear classifiers on frozen encoders across different datasets, enabling cross-dataset transfer evaluation.
@@
A widely adopted strategy in practice is to initialize model training with weights pre-trained on large-scale natural image datasets, such as ImageNet1K [20], and subsequently fine-tune the model on a medical dataset [56, 58, 68] — a paradigm known as transfer learning. Supervised ImageNet1K weights can be used either to initiate subsequent self-supervised training [56, 55], or can be directly used without further training [21]. We adopt the former approach as it better adapts the model to the fine-grained medical domain. Building upon previous research, we investigate how ImageNet1K initialization affects model generalizability across diverse datasets, while also analyzing its impact on OOD detection performance.
@@
By addressing these questions, we provide a deeper understanding of the potential of SSL in medical imaging, offering practical insights for developing more robust and generalizable models in real-world healthcare settings.
@@ 2 Related Work
Prior self-supervised benchmarks
Several studies have focused on benchmarking self-supervised strategies on natural image datasets [19, 27, 54]. However, a gap remains in the literature when it comes to benchmarking SSL methods in the medical domain using standardized datasets. Recently, Doerrich et al. [21] proposed a systematic benchmark covering convolutional and transformer-based architectures for both supervised and self-supervised strategies. Nevertheless, their approach relies heavily on ImageNet1K pre-trained encoders, without incorporating self-supervised pre-training on medical datasets. This may limit model’s ability to capture domain-specific features. Kang et al. [40] demonstrated that self-supervised pre-training outperforms supervised ImageNet1K baselines for pathology but did not explore other modalities. Huang et al. [38] analyzed SSL and semi-supervised methods with hyperparameter tuning but limited their study to four medical datasets, without assessing robustness or generalizability across diverse tasks. Our work advances SSL benchmarks in the medical domain by evaluating methods across several medical datasets, emphasizing multi-domain performance, encoder robustness, generalizability across different datasets, and adaptability to limited labeled data.
@@
Despite these advancements, there remains a notable gap: no prior work has comprehensively evaluated SSL methods for OOD detection in the medical domain, across diverse architectures and pre-training strategies. Our study aims to fill this gap by providing a benchmark that includes OOD detection experiments tailored to medical imaging.
@@
The ability of an encoder to perform effectively on datasets and domains beyond its training distribution is crucial for assessing representation quality. Li et al. [47] enhance generalizability through variational encoding with a linear-dependency regularization, while Yan et al. [73] propose a domain-generalization framework for medical image classification without domain labels in the supervised setting. Although Fedorov et al. [25] investigate the generalization of SSL methods in medical imaging, their analysis is confined to MRI data. Azizi et al. [7] propose an SSL-based representation learning method, evaluating the robustness of their approach across multiple medical imaging tasks and modalities, yet they do not compare different SSL methods. In contrast, we conduct systematic cross-dataset experiments to evaluate the true generalization capabilities of various SSL approaches.
@@
In medical domain, network initialization with pre-trained weights from large-scale datasets like ImageNet1K is a common and effective practice [7, 55, 55, 67, 68]. Tajbakhsh et al. [68] showed that CNNs pre-trained on ImageNet1K outperform those trained from scratch on medical datasets in a supervised fashion. Matsoukas et al. [55] demonstrated that SSL in the medical domain, initialized with supervised ImageNet1K weights, is effective for both ViTs and CNNs in medical imaging. To the best of our knowledge, the most comprehensive study on transfer learning in the medical domain is conducted by Taher et al. [67]. Although they emphasize domain-adapted continual pre-training, they implement it exclusively with supervised models, leaving the impact of continual pre-training on SSL largely unexplored. In our work, we investigate continual self-supervised pre-training, as well as the potential of self-supervised ImageNet1K initialization to yield more generalizable representations, an approach not previously explored in medical context.
@@ 3 Methodology
We consider the following eight discriminative SSL methods: SimCLR [14], DINO [13], BYOL [28], ReSSL [80], MoCo v3 [15], NNCLR [23], VICREG [8], and Barlow Twins [77] which are explained briefly in Appendix B. A comprehensive survey by Huang et al. [37] highlights SimCLR, MoCo, and BYOL as the most frequently adopted SSL frameworks in medical imaging research. We include the other methods to make the benchmark more comprehensive.
@@
The datasets utilized in this study are derived from MedMNIST [74], which was introduced to standardize research on medical imaging tasks. To maintain our focus on multiclass medical classification and ordinal regression, we exclude ChestMNIST from our analysis, as it only offers multi-label disease classification. As a result, our experiments span 11 MedMNIST datasets. Additional information about MedMNIST is provided in Appendix A.
@@
We employed the solo-learn library [19] with minor modifications for pre-training and linear evaluation. Our implementation details are provided in Appendix C.1, and the code will be released after acceptance. In total, we focus on five different pre-training schemes: supervised training with (1) random initialization and (2) supervised ImageNet1K initialization; self-supervised pre-training with (3) random initialization, (4) supervised ImageNet1K initialization, and (5) self-supervised ImageNet1K initialization.
@@
We evaluate various SSL methods for distinguishing between in-distribution and OOD samples, focusing on the impact of backbone architectures, pre-training strategies, and multi-domain learning on OOD detection. We use Mahalanobis Distance [46] to assign pseudo-labels by calculating the distance between input features and the nearest class-conditional Gaussian. Further details on the OOD detection metric and evaluation criteria can be found in Appendix C.5.
@@
We examine the impact of multi-domain learning by comparing models trained on single-domain MedMNIST datasets with those trained on two dataset combinations. The first combination, named Organ{A,C,S}, merges the Organ{A,C,S}MNIST datasets to represent a single-modality scenario. The second combination, referred to as Organ{A,S}PnePath, includes Organ{A,S}MNIST (CT), PathMNIST (Colon Pathology), and PneumoniaMNIST (Chest X-Ray), allowing us to assess the effect of combining different modalities. We then train SSL methods from scratch on these combined datasets to assess how dataset combination affects performance.
@@ 4.1 In-Domain Performance
We evaluate each SSL method’s in-distribution (ID) performance by pre-training a randomly initialized ResNet-50 on each dataset’s training split, then performing linear evaluation on the test split using the frozen backbone. As shown in Table 2, MoCo v3 exhibits strong performance, achieving the highest accuracy across 5 of the 11 datasets. BYOL and SimCLR also demonstrate competitive results, trailing closely behind MoCo v3. Notably, self-supervised learning outperforms supervised learning in 7 out of 11 datasets when both approaches start from random initialization.
@@
Table 2: In-domain performance of SSL methods on a randomly initialized ResNet-50 backbone, reported as Area Under the Curve (AUC) and Accuracy (ACC). Supervised learning results with random and ImageNet1K initialization are included for comparison. The highest accuracy scores achieved using SSL methods are highlighted in green while the lowest ones are highlighted in red.
Next, we examine the impact of using self-supervised ImageNet1K initialization, as opposed to random or supervised ImageNet1K initialization, on in-domain performance using MoCo v3, SimCLR, BYOL, and DINO. Table 3 presents a comparative analysis of these initialization strategies for self-supervised pre-training using ResNet-50. Although there does not appear to be a clear winner between them, self-supervised ImageNet1K initialization exhibits competitive performance relative to supervised ImageNet1K initialization, achieving the best accuracies in 5 of the 11 datasets examined. Currently, deep learning frameworks lack a standardized approach for initialization using self-supervised weights. These results can motivate developers to also include self-supervised ImageNet1K weights in deep learning libraries.
Dataset MoCo v3 SimCLR DINO BYOL
Sup. SSL Sup. SSL Sup. SSL Sup. SSL
Path 92.84 93.52 92.81 92.52 94.05 93.85 93.85 93.43
Derma 78.50 77.29 77.29 76.66 78.63 80.21 78.99 77.12
OCT 82.94 81.48 80.98 75.46 80.64 80.10 82.44 80.56
Pneumonia 92.18 94.36 94.81 93.78 93.85 92.12 93.88 91.47
Retina 50.05 51.40 51.15 52.35 46.15 55.40 54.10 41.75
Breast 88.59 84.62 87.18 90.00 88.08 89.49 88.46 85.77
Blood 97.90 98.12 97.88 97.99 98.35 98.22 98.08 98.29
Tissue 59.09 63.27 60.65 62.90 59.43 47.21 59.58 61.23
OrganA 91.85 91.70 87.69 91.02 95.13 93.43 92.70 91.58
OrganC 90.47 91.26 91.47 89.58 91.51 91.81 89.38 91.01
OrganS 79.69 77.39 77.72 78.91 78.48 79.75 80.63 76.18
@@ 4.2 Out-of-Distribution Detection
We perform several experiments to assess OOD detection performance of each SSL method. In each experiment, one dataset serves as the ID dataset, while others are treated as OOD for evaluation; resulting in 110 OOD tests ($11\times 10$) per method. The pre-trained encoder, trained on the ID dataset, is evaluated for its ability to detect OOD samples.
Figure 5 illustrates the distribution of AUROC scores for OOD detection across various SSL methods using randomly initialized ResNet-50 backbones. Among these, NNCLR and MoCo v3 achieve the highest AUROC scores, suggesting that these methods are particularly effective in learning representations that differentiate ID from OOD samples. When evaluated with a ViT-Small backbone, MoCo v3 demonstrate strong OOD detection performance, which aligns with expectations, as this method were originally developed for the ViT architecture. Detailed analysis of the effect of backbone choice on each method and in-distribution, out-of-distribution pairs $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ can be found in Appendix D.2.2.
Figure 5: AUROC score distributions for SSL methods with randomly initialized ResNet-50 backbones in OOD detection. Models are ordered by mean AUROC scores (shown above), calculated by averaging over all $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ combinations. Black dots indicate individual AUROC scores for each $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ dataset pair.
@@
Figure 6: Density distributions of AUROC scores for OOD detection, calculated across all SSL models, comparing ResNet-50 and ViT-Small with random initializations. The dashed line indicates the mean AUROC score, averaged over all $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs.
We observe that the impact of initialization on OOD detection varies depending on the choice of backbone, SSL method, and in-domain dataset distribution $\mathcal{P}_{\text{ID}}$. Generally, for smaller datasets such as BreastMNIST, RetinaMNIST, and DermaMNIST, models initialized with ImageNet1K weights demonstrate improved performance. For the ResNet-50 backbone, SSL methods such as SimCLR, ReSSL, MoCo v3, and VICReg showed improved performance in OOD detection with random initialization. In the case of ViT-Small backbone models, although differences in OOD detection scores were less pronounced, there was still a noticeable tendency for models to favor random initialization, with VICReg showing the strongest preference for random initialization among them. Detailed information about initialization preferences for each SSL method and backbone can be found in Appendix D.2.3.
Lastly, we compare OOD detection performance between models trained on multi-domain datasets (Organ{A,C,S} and Organ{A,S}PnePath) versus their constituent single-domain datasets. Figure 7 shows the mean AUROC differences across SSL methods. Our results show that for all SSL methods, Organ{A,S}PnePath consistently improves AUROC scores across most ID-OOD pairs, with the exception of a single pair. In contrast, Organ{A,C,S} generally decreases OOD detection performance across most ID-OOD pairs, particularly for models trained with MoCo v3, SimCLR, and ReSSL. These findings demonstrate that a multi-domain dataset with greater domain diversity (Organ{A,S}PnePath) provides a more substantial OOD detection performance boost compared to a more homogeneous multi-domain dataset (Organ{A,C,S}).
Figure 7: Mean average differences in OOD AUROC scores between models trained on single-domain and multi-domain datasets. The y-axis lists $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs where $\mathcal{P}_{\text{ID}}$ represents the in-domain dataset used to calculate Gaussian class centers, and $\mathcal{P}_{\text{OOD}}$ represents the OOD dataset. For clarity, only the top 6 pairs with the highest AUROC score differences and the bottom 6 pairs with the lowest differences are shown.

### v2

# Evaluating Self-Supervised Learning in Medical Imaging: A Systematic Investigation of Robustness, Generalizability, and Multi-Domain Transfer

@@ Abstract
Self-supervised learning (SSL) has emerged as a promising paradigm in medical imaging, addressing the chronic challenge of limited labeled data in healthcare settings. While SSL has shown impressive results, existing studies in the medical domain are often limited in scope, focusing on specific datasets or modalities, or evaluating only isolated aspects of model performance. This fragmented evaluation approach poses a significant challenge, as models deployed in critical medical settings must not only achieve high accuracy but also demonstrate robust performance and generalizability across diverse datasets and varying conditions. To bridge this gap, we conduct a rigorous investigation into the design space of SSL for medical imaging, evaluating 8 major SSL methods across 11 real-world medical datasets. Our analysis spans three core dimensions: (1) in-domain performance under varying label proportions (1%, 10%, and 100%), (2) cross-dataset generalization, and (3) robustness to out-of-distribution (OOD) samples. Beyond empirical evaluation, we further examine how initialization strategies, model architectures, and multi-domain pre-training contribute to SSL’s success in medical imaging. Our findings establish a principled understanding of how design choices shape SSL performance in medical imaging, providing a foundation for building more robust, generalizable, and clinically viable self-supervised models.
@@ 1 Introduction
Medical image annotation is a resource-intensive task that requires specialized domain knowledge, making it significantly more costly and laborious than annotating natural images [81, 47, 65, 40]. The scarcity of labeled medical data, coupled with the complexity of the annotation process, presents a significant challenge for building effective machine learning models in healthcare. Self-supervised learning has emerged as a powerful solution to these limitations, enabling models to learn rich representations from unlabeled data prior to task-specific fine-tuning with minimal labeled samples [25, 22, 41, 6, 2, 73, 39].
@@
While self-supervised learning (SSL) has proven effective in improving classification performance with limited labels [14, 30, 82, 13, 15], there has been limited work in evaluating the robustness and generalizability of learned models. For safe deployment in healthcare, models must perform reliably across diverse settings and recognize when to withhold predictions on out-of-distribution (OOD) samples. Although SSL methods have been evaluated for OOD detection in natural image contexts [50, 45, 37, 59, 32], their effectiveness in medical imaging remains largely under-researched. Existing studies are limited in scope: either focusing on a single imaging modality using only supervised approaches [9], examining a few SSL methods in one medical domain [11], or investigating only task-specific SSL techniques rather than general-purpose methods [12].
@@
Beyond OOD detection, an effective model should generalize across different tasks and modalities— a crucial yet underexplored aspect of SSL in medical imaging, where datasets and conditions vary widely. A model that performs well across different datasets or modalities ensures continued diagnostic support even when some imaging modalities are inaccessible, reduces the need for extensive retraining, and ensures robust performance in diverse clinical settings. Azizi et al. [7] propose an SSL-based representation learning method and assess its robustness across various medical imaging tasks and domains. However, their study focuses on a single approach, leaving it unclear whether their findings hold across different SSL paradigms.
@@
To address these gaps, we conduct a systematic investigation into the SSL design space in medical imaging, evaluating robustness and generalizability across diverse settings. We assess various SSL methods for OOD detection, comparing the effectiveness of convolutional networks (ResNet-50) with transformer-based architectures (ViT-Small). We further assess the generalizability of learned representations by training linear classifiers on frozen encoders across different datasets, enabling cross-dataset transfer evaluation.
@@
A widely adopted strategy in practice is to initialize model training with weights pre-trained on large-scale natural image datasets, such as ImageNet1K [21], and subsequently fine-tune the model on a medical dataset [58, 60, 70] — a paradigm known as transfer learning. Supervised ImageNet1K weights can be used either to initiate subsequent self-supervised training [58, 57], or can be directly used without further training [22]. We adopt the former approach as it better adapts the model to the fine-grained medical domain. Building upon previous research, we investigate how ImageNet1K initialization affects model generalizability across diverse datasets while also analyzing its impact on OOD detection performance.
@@
By addressing these questions, we provide a deeper understanding of the potential of SSL in medical imaging, establishing a principled framework for developing robust and generalizable models in real-world healthcare settings.
@@ 2 Related Work
Prior self-supervised evaluations
Several works have assessed self-supervised strategies on natural image datasets [19, 28, 56]. However, a gap remains in the literature when it comes to a systematic evaluation of SSL methods in the medical domain using standardized datasets. Recently, Doerrich et al. [22] proposed a systematic evaluation using supervised and self-supervised approaches. However, they mainly experiment with ImageNet1K pre-trained encoders without any fine-tuning on the medical data, which may limit model’s ability to capture domain-specific features. Kang et al. [41] demonstrated that self-supervised pre-training outperforms supervised ImageNet1K baselines for pathology but did not explore other modalities. Huang et al. [39] analyzed SSL and semi-supervised methods but limited their study to four medical datasets, without assessing robustness or generalizability across diverse tasks. In contrast, our work offers a comprehensive evaluation across several real-world medical datasets, focusing on multi-domain performance, encoder robustness, generalizability, and adaptability to limited labeled data.
@@
Despite these advancements, there remains a notable gap: no prior work has systematically examined SSL methods for OOD detection in the medical domain, across diverse architectures and pre-training strategies. Our study aims to fill this gap by conducting a rigorous analysis of OOD detection across diverse medical imaging conditions.
@@
The ability of an encoder to perform effectively on datasets and domains beyond its training distribution is crucial for assessing representation quality. Li et al. [49] enhance generalizability through variational encoding with a linear-dependency regularization, while Yan et al. [75] propose a domain-generalization framework for medical image classification without domain labels in the supervised setting. Although Fedorov et al. [26] investigate the generalization of SSL methods in medical imaging, their analysis is confined to MRI data. Azizi et al. [7] propose an SSL-based representation learning method, evaluating the robustness of their approach across multiple medical imaging tasks and modalities, yet they do not compare different SSL methods. In contrast, we conduct systematic cross-dataset experiments to evaluate the true generalization capabilities of several SSL approaches.
@@
In medical domain, network initialization with pre-trained weights from large-scale datasets like ImageNet1K is a common and effective practice [7, 57, 57, 69, 70]. Matsoukas et al. [57] demonstrated that SSL in medical domain, initialized with supervised ImageNet1K weights, is effective for both ViTs and CNNs in medical imaging. To the best of our knowledge, the most comprehensive study on transfer learning in medical domain is conducted by Taher et al. [69]. Although they emphasize domain-adapted continual pre-training, they implement it exclusively with supervised models, leaving the impact of continual pre-training on SSL largely unexplored. In our work, we investigate continual self-supervised pre-training, and the potential of self-supervised ImageNet1K initialization to yield better representations for in-domain performance and OOD detection, an approach not previously explored in medical context.
@@ 3 Methodology
We consider the following eight discriminative SSL methods: SimCLR [14], DINO [13], BYOL [29], ReSSL [82], MoCo v3 [15], NNCLR [24], VICREG [8], and Barlow Twins [79] which are explained briefly in Appendix B. A comprehensive survey by Huang et al. [38] highlights SimCLR, MoCo, and BYOL as the most frequently adopted SSL frameworks in medical imaging research. We include the other methods to make our study more comprehensive.
@@
We used real-world medical datasets, including NCT-CRC-HE-100K, CRC-VAL-HE-7K [43, 42], HAM10000 [71], OCT dataset [44], pediatric chest X-Ray images [44], DeepDRiD Retina dataset [52], breast ultrasound images [4], microscopic peripheral blood cell images [1], BBBC051 [54] and LiTS [10], all curated and standardized under the MedMNIST framework [76]. For the LiTS dataset, we followed the MedMNIST adaptation, which transforms the original segmentation task into an organ classification problem by extracting regions of interest.
To maintain our focus on multiclass medical classification and ordinal regression, we exclude ChestMNIST from our analysis, as it only offers multi-label disease classification. Our experiments span 11 MedMNIST datasets, additional information being provided in Appendix A.
@@
We employed the solo-learn library [19] with some modifications for pre-training, OOD detection and linear evaluation. Our implementation details are provided in Appendix C.1, and the code will be released after acceptance. In total, we focus on five different pre-training schemes: supervised training with (1) random initialization and (2) supervised ImageNet1K initialization; self-supervised pre-training with (3) random initialization, (4) supervised ImageNet1K initialization, and (5) self-supervised ImageNet1K initialization.
@@
We evaluate various SSL methods for distinguishing between ID and OOD samples, focusing on backbone architectures, pre-training strategies, and multi-domain learning. We follow the approach defined in [48], where feature representations $\mathbf{f}(x)\in\mathbb{R}^{d}$ are assumed to follow a class-conditional multivariate Gaussian, with parameters estimated from the ID training data $\mathcal{P}_{\text{ID}}$ (i.e., samples used to train the encoder). For each class $c$, the Mahalanobis distance is computed as
and the confidence score is given by
$S(x)=-\min_{c}D_{\text{M}}(x,c).$
Samples are considered OOD if they originate from a different dataset or subset (denoted as $\mathcal{P}_{\text{OOD}}$) than the one used for training. OOD detection performance is evaluated using threshold-independent metrics such as AUROC and AUPR. Further details are provided in Appendix.
@@
We examine the impact of multi-domain learning by comparing models trained on single-domain MedMNIST datasets with those trained on two dataset combinations. The first combination, named Organ{A,C,S} merges the Organ{A,C,S}MNIST datasets to represent a single-modality scenario. The second combination, referred to as Organ{A,S}PnePath, includes Organ{A,S}MNIST (CT), PathMNIST (Colon Pathology), and PneumoniaMNIST (Chest X-Ray), allowing us to assess the effect of combining different modalities. We then train SSL methods from scratch on these combined datasets to assess how dataset combination affects performance.
@@ 4.1 In-Domain Performance
We assess each SSL method’s in-distribution performance by pre-training ResNet-50 on the train set and evaluating with a linear classifier on the test set using frozen features. MoCo v3 exhibits strong performance, achieving the highest accuracy across 5 of the 11 datasets. BYOL and SimCLR also demonstrate competitive results, trailing closely behind MoCo v3. Notably, self-supervised learning outperforms supervised learning in 7 out of 11 datasets when both approaches start from random initialization. Numerical results can be found in Appendix D.1.
@@
[…]
@@ 4.2 Out-of-Distribution Detection
We perform several experiments to assess OOD detection performance of each SSL method. In each experiment, one dataset serves as the ID dataset, while others are treated as OOD for evaluation; resulting in 110 OOD tests (11×1011$\times$ 1011 × 10) per method. The pre-trained encoder, trained on the ID dataset, is evaluated for its ability to detect OOD samples.
Figure 5: AUROC score distributions for models with randomly initialized weights and a ResNet-50 backbone in OOD detection. Models are ordered by mean AUROC scores (shown above), calculated by averaging over all $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ combinations. Black dots indicate individual AUROC scores.
Figure 5 illustrates the distribution of AUROC scores for OOD detection across various SSL methods using randomly initialized ResNet-50 backbones. Among these, NNCLR and MoCo v3 achieve the highest AUROC scores, suggesting that these methods are particularly effective in learning representations that differentiate ID from OOD samples. When evaluated with a ViT-Small backbone, MoCo v3 demonstrated strong OOD detection performance, which aligns with expectations as it was originally developed for the ViT architecture. Detailed analysis of effect of backbone choice on each method and in-distribution, OOD pairs $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ can be found in Appendix D.2.2.
@@
Figure 6: Mean AUROC score differences between ResNet-50 and ViT-Small models calculated over all $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs. Negative values indicate a performance advantage for ViT-Small, while positive values favor ResNet-50. All differences are statistically significant ($p<0.05$), demonstrating a clear overall advantage of ViT-Small models for OOD detection.
We observe that the impact of initialization on OOD detection varies depending on the choice of backbone, SSL method, and in-domain dataset distribution $\mathcal{P}_{\text{ID}}$. Generally, for smaller datasets such as BreastMNIST, RetinaMNIST, and DermaMNIST, models initialized with ImageNet1K weights demonstrate improved performance. For the ResNet-50 backbone, SSL methods such as SimCLR, ReSSL, MoCo v3, and VICReg enhanced OOD detection performance with random initialization. In the case of ViT-Small backbone models, although differences in OOD detection scores were less pronounced, there was still a noticeable tendency for models to favor random initialization, with VICReg showing the strongest preference.
Moreover, initializing the models with self-supervised trained ImageNet1K weights led to significant differences in AUROC scores compared to starting with supervised ImageNet1K weights for several datasets. In particular, significant negative differences (i.e., lower AUROC with self-supervised initialization) were observed for BreastMNIST ($\Delta\mathrm{AUC}=-0.12$, 95% CI: ±0.18$\pm$ 0.18± 0.18), PathMNIST ($\Delta\mathrm{AUC}=-0.09$, 95% CI: ±0.13$\pm$ 0.13± 0.13), and RetinaMNIST ($\Delta\mathrm{AUC}=-0.03$, 95% CI: ±0.09$\pm$ 0.09± 0.09). Conversely, significant positive differences (i.e., higher AUROC with self-supervised initialization) were found for PneumoniaMNIST ($\Delta\mathrm{AUC}=0.08$, 95% CI: ±0.15$\pm$ 0.15± 0.15) and TissueMNIST ($\Delta\mathrm{AUC}=0.08$, 95% CI: ±0.19$\pm$ 0.19± 0.19). Detailed information about initialization preferences for each SSL method and backbone can be found in Appendix D.2.3.
Figure 7: Mean average differences in AUROC scores between models trained on multi-domain and single-domain datasets. The y-axis lists $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs where. Positive values (red) indicates a favor for models trained on multi-domain dataset compared to its single-domain counterpart. For clarity, only the top/bottom 6 pairs with the highest/lowest AUROC score differences are shown.
Lastly, we compare OOD detection performance between models trained on multi-domain datasets (Organ{A,C,S} and Organ{A,S}PnePath) versus their constituent single-domain datasets. Figure 7 shows that for all SSL methods, Organ{A,S}PnePath consistently improves AUROC scores across most $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs, with the exception of a single pair. In contrast, Organ{A,C,S} generally decreases OOD detection performance across most $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs, particularly for models trained with MoCo v3, SimCLR, and ReSSL. These findings demonstrate that a multi-domain dataset with greater domain diversity (Organ{A,S}PnePath) provides a more substantial OOD detection performance boost compared to a more homogeneous multi-domain dataset (Organ{A,C,S}).